### v1

# Hadamax Encoding: Elevating Performance in Model-Free Atari

Abstract
Neural network architectures have a large impact in machine learning. In reinforcement learning, network architectures have remained notably simple, as changes often lead to small gains in performance. This work introduces a novel encoder architecture for pixel-based model-free reinforcement learning. The Hadamax (Hadamard max-pooling) encoder achieves state-of-the-art performance by max-pooling Hadamard products between GELU-activated parallel hidden layers. Based on the recent PQN algorithm, the Hadamax encoder achieves state-of-the-art model-free performance in the Atari-57 benchmark. Specifically, without applying any algorithmic hyperparameter modifications, Hadamax-PQN achieves an 80% performance gain over vanilla PQN and significantly surpasses Rainbow-DQN. For reproducibility, the full code is available on GitHub.
[…]
1 Introduction
Figure 1: Performance versus GPU hours in the full Atari-57 domain at 200M environment frames. The application of our Hadamard max-pooling encoder on PQN yields significant performance improvements over a current state-of-the-art model-free method, Rainbow, while remaining more than an order of magnitude faster.
Ever since reinforcement learning (RL) algorithms [51] surpassed human players on the Atari-57 benchmark [6, 36, 37], progress has been driven mainly by various algorithmic innovations [15, 48].
Compared with the field of supervised learning (SL), the deep learning components of RL have remained relatively simple, usually consisting of a few convolutional layers (for image-based tasks) followed by fully connected layers [37, 27]. So far, the most common encoder modification in image-based RL tasks has been the integration of a ResNet encoder [13], inspired by its wide use in supervised learning architectures [25]. Several further approaches have been explored to scale the deep learning architecture, but findings indicate that scaling pixel-based RL remains a significant challenge [40, 41], and finds greater success in either low-dimensional state-based continuous control [34, 24] or complex model-based architectures [45, 23].
In this work, we revisit the assumption that modifications to baseline network architectures correspond with instabilities in RL. We build on top of the recent Parallelised Q-Network (PQN), which reinvented DQN to function without the replay buffer and target network, while profoundly increasing performance [17]. This is done by combining recent advances in Hadamard representations [30] with max-pooling found in the ResNet encoder structures [25, 13]. Specifically, we augment the state-of-the-art PQN algorithm with a Hadamard-maxpooling (Hadamax) encoder. This paper’s contributions can be summarized as follows:
  • A novel deep learning architecture is proposed to improve usual pixel-based convolutional encoder architectures for model-free RL. This design shows an alternative direction of encoder synthesis in RL, as compared to the widely used deeper ResNet architectures.
  • Without applying any algorithmic or hyperparameter modifications, Hadamax-PQN achieves an 80% performance gain in the full Atari-57 suite over the recent PQN baseline [17]. Additionally, the proposed encoder changes allow Hadamax-PQN to significantly surpass Rainbow-DQN [27] while remaining more than an order of magnitude faster, setting a stronger baseline for model-free RL on Atari.
2 Related Work
Different neural network architectures are applied in RL to enhance the performance in online settings [36, 5, 13, 27, 17, 34] as well as offline limited data settings [7, 47, 10, 48, 38]. In this work, we focus on agents in the high-dimensional Atari-57 domain [6], a diverse and commonly-used challenging benchmark with discrete actions and pixel-based input.
Network development in RL for Atari: Deep Q-learning (DQN) [36, 37] achieves human-level performance on Atari games for the first time in the RL history by using three convolution layers (with ReLU) followed by fully-connected layers. Due to its simplicity and efficiency, this classic architecture is used for many later works, such as Double DQN (DDQN) [52], Dueling DQN [55], Noisy DQN [14], Categorical DQN (C51) [5] and Rainbow-DQN  [27]. The recent Parallel Q-learning (PQN) [17] algorithmically simplifies DQN and uses LayerNorm [4] to provably stabilize optimization. C51 [5] and R2D2 [29] enhance the output layer using categorical distributions and a recurrent network, respectively. In the context of model-based RL, Recurrent State-Space Models (RSSM)  [20, 21, 23], image augmentation [33], forward prediction [47, 39], residual architectures [13, 44] and transformers [1] have also been explored to solve Atari. Impala [13] introduces a deeper ResNet-15 encoder structure with 6 residual blocks, allowing for high data efficiency under distributional training. BBF [48] further widens the Impala encoder, achieving state-of-the-art performance on the Atari-100k benchmark. SPR [47], using DQN’s architecture with a self-prediction objective, also improves data efficiency. For model-based methods, residual architectures [57, 54], transformers [58] and diffusion models [3] are being increasingly leveraged to boost sample efficiency. Our work focuses on model-free agents in the Atari-57 benchmark, where relatively modest algorithmic architectures are used, and a large amount of environment interactions is allowed.
Vectorized RL: Since the development of JAX [9], parallel and vectorized training of reinforcement learning (RL) agents has become a promising area of research, offering significant performance and scalability improvements. Physics simulation engines and tools that are compatible with JAX have emerged to support this paradigm, including Brax [16], a physics simulation engine optimized for high-speed differentiable environments; Gymnax [32], a lightweight, JAX-based version of classic Gym environments; Jumanji [8], a suite of combinatorial and decision-making environments tailored for JAX; and EnvPool [56], a high-throughput environment execution engine with up to 20x speedup compared to Python. To complement these environments, a growing ecosystem of reinforcement learning libraries built entirely in JAX has been developed. PureJaxRL [35] implements standard RL algorithms entirely end-to-end in JAX, enabling parallel execution across thousands of environments. JaxMARL [43] focuses on multi-agent reinforcement learning, demonstrating strong acceleration of existing algorithms. Additionally, cleanrl [28], a library providing high-quality and reproducible RL implementations, also includes several JAX-based implementations. Our work builds upon PQN, which leverages EnvPool and PureJaxRL, achieving greater efficiency compared to conventional PyTorch-based implementations. With the Hadamax encoder, we further architecturally improve PQN to the point that it significantly surpasses Rainbow-DQN, while remaining more than an order of magnitude faster.
3 Preliminaries
As a background, we briefly explain general value-based RL and the recent PQN algorithm, which is extended with our proposed encoder.
3.1 Reinforcement Learning and Value-based Methods
We consider a Markov Decision Process (MDP), defined by the tuple $<\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma>$, with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $\mathcal{P}$, reward function $\mathcal{R}$ and discount factor $\gamma\in[0,1)$. An agent in state $s_{t}\in\mathcal{S}$ at timestep $t$, taking action $a_{t}\in\mathcal{A}$ observes a reward $r_{t}\sim\mathcal{R}(s_{t},a_{t})$ and next state $s_{t+1}\sim\mathcal{P}(s_{t},a_{t})$. The goal is to learn an optimal policy $\pi^{*}:\mathcal{S}\rightarrow\mathcal{A}$ that can maxmize the expected return $G(s_{t})=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}\mid s_{t}=s\right]$ over all possible trajectories. Unlike policy-based or actor-critic methods [46, 19] that optimize the policy, value-based methods [36] learn a state-action value function $Q(s,a)$. Once the optimal Q-function is learned, the optimal policy is implicitly defined by selecting greedy actions $\pi^{*}(s)=\text{argmax}_{a}Q^{*}(s,a)$. Q-learning is the most widely used value-based algorithm. It learns $Q(s,a)$ through temporal difference (TD) learning. The update rule is:
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha[r+\gamma\text{max}_{a^{\prime}\in\mathcal{A}}Q(s_{t+1},a^{\prime})-Q(s_{t},a_{t})],$ (1)
where $\alpha$ is the learning rate. Over time, this iterative process allows the Q-function to converge to the optimal value function $Q^{*}(s,a)$, from which the optimal policy can be derived.
Deep Q-Network (DQN) [36] extends Q-learning by using a deep neural network to approximate the Q-function. The network is trained to minimize the difference between the predicted Q-values and the target values, typically using a loss function such as mean squared error:
$\mathcal{L}(\theta)=\mathbb{E}_{s_{t},a_{t},r_{t},s_{t+1}\sim D}[(r_{t}+\gamma\text{max}_{a^{\prime}\in\mathcal{A}}Q(s_{t+1},a^{\prime};\theta^{-})-Q(s_{t},a_{t};\theta))^{2}]$ (2)
where $\theta$ and $\theta^{-}$ are the parameters of the Q-network and are the parameters of a target network that is periodically updated. $D$ is the experience replay buffer from which mini-batches are sampled.
3.2 Parallelised Q-Network
PQN is a simplified deep online Q-learning algorithm. By parallelizing vectorized environments and normalizing neural networks (LayerNorm), PQN can stabilize the training even without a target network and replay buffer. Moreover, it is compatible with pure-GPU training, leading to efficient training on Atari tasks. More specifically, PQN makes the following modifications compared to the original DQN:
$\lambda$-return: Unlike the original DQN uses 1-step return, PQN leverages a more stable $\lambda$-return. The loss in Equation 2 thus becomes:
$\mathcal{L}(\theta)=\mathbb{E}_{\text{trajs}}[(r_{t}+\gamma(\lambda G_{t+1}^{\lambda}+(1-\lambda)\text{max}_{a^{\prime}\in\mathcal{A}}Q(s_{t+1},a^{\prime};\theta))-Q(s_{t},a_{t};\theta))^{2}],$ (3)
where $G^{\lambda}$ is the $\lambda$-return. When $\lambda=0$ it will be similar to Q-learning, and if $\lambda=1$ it is equivalent to the Monte Carlo update, which uses the full return until the episode ends.
LayerNorm: PQN adds LayerNorm for the output of convolution / MLP layers before the ReLU activation functions, which helps stabilize the training process.
Vectorized Environments: Vectorization enables fast collection of parallel trajectories from independent environments. It enhances exploration and stabilizes the training [17].
Removal of replay buffer and target network: Since the whole training process happens on GPU, removing the replay buffer can largely reduce memory and thereby accelerate training. As a result of the training stability, the target network is also eliminated.
4 Hadamax Encoder
Figure 2: Encoder architectures of DQN, PQN , the proposed Hadamard max-pooling (Hadamax) encoder and the Impala ResNet-15 encoder (from left to right). In the Hadamax encoder, down-sampling is facilitated by max-pooling operators. Furthermore, we apply a Hadamard product between parallel representation layers. The implementation is straightforward and can be found in Appendix B. These changes allow for a substantial increase in algorithm performance, while keeping general encoder structure, convolutional depth and algorithmic hyperparameters unchanged.
The first human-level performance in the Atari-57 domain was achieved with the ’Nature’ DQN encoder design [37]. The general effectiveness of this architecture, as well as the problems with scaling in deep RL, has led to this architecture’s use even in the modern state-of-the-art algorithms such as PQN [17]. In this section, we provide the reasoning and implementation of the proposed Hadamard max-pooling augmentation of the original DQN encoder. For reproducibility purposes, we refer the reader to a detailed implementation of the proposed architecture in Appendix B.
4.1 Design Choice 1: Down-sampling by Max-pooling
As pixel-based observations are high-dimensional, the encoder must effectively compress the state representation to enable the downstream RL algorithm to converge within a reasonable number of updates. In the conventional DQN encoder, this compression is achieved by the convolutional operations (See Fig. 2), where the compression is determined by the convolutional kernel size and stride. In contrast, when examining the well-known and widely used Impala ResNet-15 encoder in RL [13], max-pooling is responsible for the bulk of feature compression. The resulting effect is that minimizing convolutional strides and adding max-pooling allows for the selection of a more dense representation of convolutional features, and subsequently emphasizes the strongest signals. Additionally, the use of max-pooling adds a slight translation invariance to the important features. We therefore hypothesize that the use of max-pooling in RL is, although widely implemented in supervised learning, relatively overlooked. In the Hadamax encoder, convolutional down-sampling is therefore replaced by max-pooling operators. Furthermore, in contrast to the average-pooling used by the original supervised learning ResNet architecture [25], the Hadamax encoder max-pools the final features before flattening to the linear layer. Since value functions in RL should be able to show strong correlations with the most important features, average-pooling before the linear layer will achieve the opposite, as it smoothens out feature importance.
The max-pooling design choices; max-pooling and downsampling instead of convolutional down-sampling, followed by max-pooling without down-sampling before flattening, are thus respectively influenced by the ResNet-15 (Impala) RL encoder and the original ResNet. However, in stark contrast to both residual encoders mentioned, the Hadamax encoder remains shallow (3 convolutional layers), and therefore no residual connections need to be applied.
4.2 Design Choice 2: Application of Hadamard Representations
Although multiplicative interactions have been commonly used in Deep Learning architectures [50, 53, 11], their application in RL remains limited. Recent work however has shown that the effective rank (ER [31, 18]) and downstream performance improved when training deep RL in the Atari domain, by defining hidden layers as Hadamard products [30]. Hadamard products between hidden layers enable richer high-dimensional interactions within the representation space, without increasing hidden layer dimensionality. This leads to more network capacity without explicitly scaling the network, which is often unstable in RL. Specifically, any hidden layer $z^{j}\in\mathcal{Z}$, with layer depth $j$, will be the Hadamard product of two parallel layers connected to the preceding hidden layer $z^{j-1}$:
$z^{j}=f(z^{j-1}A^{j-1}_{1})\odot f(z^{j-1}A^{j-1}_{2}),$ (4)
where $A$ is a weight matrix, $f(*)$ is a nonlinear activation and the bias layers are left out for simplicity. As PQN employs layer normalization for training stability, and every representation is max-pooled, the final Hadamax representation layers can be defined as:
$z^{j}=MP\bigg{(}f\big{(}LN(z^{j-1}A^{j-1}_{1})\big{)}\odot f\big{(}LN(z^{j-1}A^{j-1}_{2})\big{)}\bigg{)}.$ (5)
Where LN and MP represent layer normalization and max-pooling, respectively. It is worth noting that contrary to recent work on Hadamard representations [30], we show the possibility of successfully applying Hadamard products to zero-saturating activation functions such as ReLU or GELU [26]. We believe this is possible due to the relative training stability increase of PQN over DQN, as a result of applying LayerNorm and the removal of the target network and replay buffer. This training stability correlates with a minimal amount of dead neurons in the representation [49], which even leads to the ability to do element-wise multiplication of zero-saturating (sparse) neurons without increasing dead neurons.
4.3 Design Choice 3: Gaussian Error Linear Unit
The Gaussian Error Linear Unit (GELU) is used in various neural network architectures, the most notable applications being in transformer-based architectures such as BERT [12] and GPT [42].
It is defined as:
$\text{GELU}(x)=x\Phi(x)$
Figure 3: ReLU and GELU.
where $\Phi(x)$ is the cumulative distribution function of the standard normal distribution. Equivalently, it can be expressed using the error function as:
$\text{GELU}(x)=0.5x\left(1+\text{erf}\left(\frac{x}{\sqrt{2}}\right)\right)$
In contrast to the ReLU, which converts negative inputs to zero, GELU permits small negative values to pass through in a softened form (See Fig. 3), allowing more stable gradient flow for negative inputs. Overall, GELU has been shown to improve performance in various deep learning tasks, including computer vision and natural language processing [26]. In the Hadamax encoder, we therefore replace all the original ReLU activation functions with the GELU.
5 Experiments
We compare our agent against widely used model-free RL baselines across 57 Atari games. Through experiments, we aim to answer: (i) do agents equipped with Hadamax encoders outperform those using conventional encoders? (ii) what are the reasons behind Hadamax’s superior performance? (iii) what is the impact of each proposed design choice?
Baselines: We compare our method with the following baselines: (1) DQN [36], a pioneer RL method that uses a deep neural network to play Atari, achieving human performance. (2) C51 [5], Rainbow [27], a state-of-the-art model-free method, combining various algorithmic and architectural techniques together. (3) PQN [17], a recent novel parallel Q-learning network without a replay buffer and target network. In terms of performance, PQN is on par with C51, while remaining algorithmically less complex than DQN. Our final baseline is (4) PQN (ResNet-15), which combines PQN with the more complex Impala CNN architecture, used throughout modern state-of-the-art RL algorithms as a drop-in replacement for the conventional Nature encoder [13, 48].
Figure 4: The Atari-57 domain.
Environments: The full 57-game Atari domain [6] is used as a standardized benchmark for evaluating our algorithm’s performance. In line with best practices in the field, we focus on the median human-normalized score over all 57 games [37, 27, 21, 17]. To manage computational load, ablations are done on 40M frames, while comparison with baselines is done at the official 200M frame scores. Note that there can be relative differences between performances in 40M and 200M frames, as the epsilon-greedy coefficient $\epsilon$ is scaled down over the total training time. An algorithm seed initialized to run for 40M frames will therefore have a different convergence curve towards 40M than the same algorithm initialized for a 200M frames seed. We refer the reader to more detailed descriptions of environments and implementations of baseline agents in Section C.3.
5.1 Hadamax-PQN: Results
The full 200M frame training curves for PQN, PQN (ResNet-15) and Hadamax-PQN are shown in Fig. 5 (left). The Hadamax encoder clearly yields benefits over the widely used Impala ResNet-15 encoder [13], and causes PQN to significantly surpass Rainbow-DQN [27]. Although the original paper shows that PQN is able to beat Rainbow-DQN when training for around 260M environment frames [17], Hadamax-PQN reaches this score at around 90M frames. Another commonly used scoring method, the Atari-57 score profile, can be seen in Fig. 5 (right). Note that the scores used in this research for DDQN, C51 and Rainbow have been taken from the original papers, and are generally higher than their practical implementations on various GitHub repositories. For details on how to compute the median human-normalized score and the Atari score profile, we refer the reader to Appendix D.
Figure 5: Median Human-Normalized performance training PQN, PQN (Resnet-15) and Hadamax-PQN in the Atari domain over 57 games, 200M frames and 5 seeds (left), and the Atari-57 score profile (right). The Atari-57 score profile illustrates the percentage of games that exceed the normalized score threshold on the x-axis.
The effect of the Hadamax encoder on the baseline PQN on a per-game basis can be seen in Fig. 6. The results show a significant performance increase over the baseline, with over 17 games having more than 100% improvement, compared to only one single game having more than a 50% decrease in performance. The per-game improvements over the Rainbow-DQN baseline can be seen in Appendix E.3. For each individual game’s training curve and the final 200M frame score table, we refer the reader to Appendix F. To the best of our knowledge, the implementation of the Hadamax encoder is one of the biggest recorded non-algorithmic improvement over a recent competitive RL baseline, and it does not involve any complex hyper-parameter tuning.
Figure 6: Per-game improvement of Hadamax-PQN over PQN (Log Scale).
5.2 Does Hadamax Generalize Beyond PQN?
Figure 7: C51 with and without a Hadamax encoder on Atari-10.
The Hadamax encoder not only enhances the performance of PQN, but also works effectively with other reinforcement learning agents. To showcase this, the C51 algorithm is evaluated on the Atari-10 benchmark for 40M environment frames. As shown in Figure 7, a direct implementation of the Hadamax encoder to the C51 algorithm boosts the performance by approximately $70\%$ on Atari-10 [2]. Similar to PQN, the algorithmic hyperparameters for Hadamax-C51 remain exactly the same as for the C51 baseline from cleanrl [28]. These improvements suggest that the Hadamax encoder is able to be implemented as a strong default encoder for multiple algorithms in the Atari domain. For more information on the Atari-10 benchmark and the corresponding score normalization metrics, we refer the reader to Section D.3.
5.3 Effective Rank and Dead Neurons
In order to obtain clues about the stabilizing effects of the proposed Hadamax encoder, the effective rank of the hidden layers is investigated during training [31, 18], as well as the amount of dead neurons [49].
Figure 8: Effective rank [31] of the 4 hidden layers for both the baseline PQN and the Hadamax-PQN setting. Although there is no visible difference between the first and final layer, the deeper convolutional layers show a lower effective rank in the baseline setting, as well as a stronger rank decay during training.
The effective rank of a feature matrix for a threshold $\delta$ ($\delta=0.01$), denoted as $srank_{\delta}(\Phi)$, is given by $srank_{\delta}(\Phi)=\min\left\{k:\frac{\sum_{i=1}^{k}\sigma_{i}(\Phi)}{\sum_{i=1}^{d}\sigma_{i}(\Phi)}\geq 1-\delta\right\}$, where $\{\sigma_{i}(\Phi)\}$ are the singular values of $\Phi$ in decreasing order, i.e., $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$. The effective rank portrays a measure of network capacity i.e. the amount of information that can be approximated in a certain hidden layer.
We investigate the differences in effective rank between the baseline PQN and Hadamax-PQN. To find clues for Hadamax’s strong improvements on certain environments, the differences are visualized on a random subset of 5 high-improvement environments from Fig. 6. The effective rank of the encoder’s representation layers while training for 200M frames can be seen in Fig. 8. The plots show that there are minimal differences in effective rank in the first and last hidden layer of the encoder. However, in the baseline PQN encoder, the deeper convolutional layers show a more prominent decay in rank during training, as well as a reduced initial effective rank. As mentioned in Section 4, the increase in effective rank in the deeper convolutional layers can largely be credited to the use of Hadamard representations.
Figure 9: Fraction of dead neurons over 200M frames.
Further investigation into the penultimate layer’s fraction of dead neurons shows a small decrease from the baseline (see Fig. 9). The percentage of dead neurons in the final hidden layer is calculated by finding neurons that have a variance of less than $10^{-4}$ over the batch dimension. In practice, this metric generalizes well to any activation function (ReLU, GELU, Tanh). After training for 200M frames, both the baseline PQN and Hadamax-PQN have less than 8% dead neurons, which remains extremely low compared to DQN [49]. We therefore do not expect a substantial correlation between the small reduction in dead neurons and the performance. However, in contrast to recent work on Hadamard representations [30], who showed that the DQN algorithm exhibits instability when multiplying ReLU-activated neurons, we show that it is possible to use Hadamard products on zero-saturating activations. We believe the inherent stability of the PQN algorithm and its corresponding low fraction of dead neurons allows for successful Hadamard multiplication of linear-unit activations like ReLU or GELU.
5.4 Which Design Choice is most Important?
As described in Section 4, the Hadamax encoder differs from PQN’s conventional Nature CNN encoder in three areas: (1) applying max-pooling (2) using Hadamard representations and (3) GELU-activated hidden layers. The precise influence of each component of the Hadamax encoder remains to be determined. An ablation analysis over these areas is therefore done on 40M environment frames in the full Atari-57 suite. The ablations are defined as implementation subtractions from the original Hadamax architecture in Fig. 2. The result of the ablation study is shown in Fig. 10(a). Next to the ablations, the effects of direct additions of our design choices on the baseline PQN are investigated. The results of the addition analysis can be seen in Fig. 10(b).
(a) Hadamax ablations.
(b) Naive PQN additions.
Figure 10: Ablations of Hadamax-PQN, each represented as a subtraction from the full Hadamax architecture (a), and naive architectural additions to the baseline PQN (b).
Over a training period of 40M frames, the subtraction of max-pooling leads to the largest decay in performance. Note that when max-pooling is subtracted from our architecture, we return the convolutional strides to their original values, in order to still retain feature compression. The importance of down-sampling with max-pooling strengthens our hypothesis that a selection of the most prominent features is key when working with high-dimensional observation spaces in the Atari domain. The use of convolutional Hadamard representations is also an important component, showing that the increase in effective rank paired with other benefits such as high-order interactions [11], have a strong correlation with downstream performance. Finally, the GELU activation has the lowest importance, although its contribution as compared to the ReLU still remains substantial for such a small architectural component. Notably, if the ablations are compared to the effects of directly implementing a single design choice on the baseline (see Fig. 10(b)), it becomes clear that the overall combination of all three components is a key factor. For an experimental analysis with two deeper Hadamax encoders, we refer the reader to Section E.1.
6 Conclusions and Future Work
This paper introduced the Hadamax encoder architecture, augmenting the conventional pixel-based Nature CNN architecture with Hadamard representations, while down-sampling using max-pooling instead of convolutional strides. Furthermore, the Gaussian Error Linear Unit activation was implemented to improve training stability. The application of these fundamental changes to PQN’s encoder, while preserving its original structure, allowed for a profound increase in performance over several model-free baselines. Specifically, we reach an almost two-fold performance gain over PQN, and surpass Rainbow-DQN’s official 200M frame score after just 90M frames, while remaining an order of magnitude faster. Additional results on C51 show that the Hadamax encoder remains effective across algorithms.
Due to computational constraints, limitations of our work include the limited testing of the Hadamax encoder on the more complex algorithms such as BBF [48] on the Atari-100k benchmark or on a state-of-the-art model-based algorithm such as Dreamer [22]. However, as seen by the performance improvement on C51 in Fig. 7, we expect a certain degree of generalization. Another limitation is that the Hadamax encoder, due to its higher complexity, accounts for some extra computational overhead compared to PQN’s conventional Nature CNN architecture, as seen in Fig. 1.
This paper takes an important step forward in encoder synthesis for RL, discovering an alternative for the more-often used deeper ResNet architectures to optimize performance. An interesting avenue for future work is to investigate scaling of the Hadamax encoder, as it already achieves significant performance improvements using only 3 convolutional layers. Finding ways to scale Hadamax in either width or depth could yield even stronger improvements. Another interesting direction would include looking at integrating a MoE-style prediction head in the Hadamax encoder, since MoE does not affect the base encoder [41]. Furthermore, as Hadamax-PQN does not come with any algorithmic or hyperparameter changes, it can be used as a new baseline to build algorithmic improvements upon. Specifically, since hard-exploration games are generally not suited for PQN’s epsilon-greedy exploration regime, augmenting PQN-Hadamax with novel exploration techniques might further bridge the gap in performance between model-free and complex model-based algorithms such as DreamerV3 or Muzero [23, 45].
Appendix
[…]
Appendix A Impact Statement
This work shows that architectural innovations like the Hadamax encoder can drive significant progress in reinforcement learning. By enabling more efficient and accessible AI, it encourages broader adoption and exploration of learning systems across diverse real-world domains.
Appendix B Hadamax Encoder Code
We provide the full JAX-based code of the Hadamax encoder for reproducibility purposes.
1# Input = input_obs, a frame-stacked Atari observation
2x = jnp.transpose(input_obs, (0, 2, 3, 1))
3x = x / 255.0
4# First block
5x1 = nn.Conv(32, kernel_size=(8, 8), strides=(1, 1), padding="SAME",
6 kernel_init=nn.initializers.xavier_normal())(x)
7x2 = nn.Conv(32, kernel_size=(8, 8), strides=(1, 1), padding="SAME",
8 kernel_init=nn.initializers.xavier_normal())(x)
9x1 = normalize(x1) # Normalize before activation
10x2 = normalize(x2) # Normalize before activation
11x1 = nn.gelu(x1) # Apply activation
12x2 = nn.gelu(x2) # Apply activation
13x = x1 * x2 # Hadamard product
14x = max_pool(x, window_shape=(4, 4), strides=(4, 4), padding="SAME")
15# Second block
16x1 = nn.Conv(64, kernel_size=(4, 4), strides=(1, 1), padding="SAME",
17 kernel_init=nn.initializers.xavier_normal())(x)
18x2 = nn.Conv(64, kernel_size=(4, 4), strides=(1, 1), padding="SAME",
19 kernel_init=nn.initializers.xavier_normal())(x)
20x1 = normalize(x1) # Normalize before activation
21x2 = normalize(x2) # Normalize before activation
22x1 = nn.gelu(x1) # Apply activation
23x2 = nn.gelu(x2) # Apply activation
24x = x1 * x2 # Hadamard product
25x = max_pool(x, window_shape=(2, 2), strides=(2, 2), padding="SAME")
26# Third block
27x1 = nn.Conv(64, kernel_size=(3, 3), strides=(1, 1), padding="SAME",
28 kernel_init=nn.initializers.xavier_normal())(x)
29x2 = nn.Conv(64, kernel_size=(3, 3), strides=(1, 1), padding="SAME",
30 kernel_init=nn.initializers.xavier_normal())(x)
31x1 = normalize(x1) # Normalize before activation
32x2 = normalize(x2) # Normalize before activation
33x1 = nn.gelu(x1) # Apply activation
34x2 = nn.gelu(x2) # Apply activation
35x = x1 * x2 # Hadamard product
36x = max_pool(x, window_shape=(3, 3), strides=(1, 1), padding="SAME")
37# Flatten for MLP layer
38x = x.reshape((x.shape[0], -1))
39x = nn.Dense(512, kernel_init=nn.initializers.he_normal())(x)
40x = normalize(x)
41x = nn.gelu(x)
42x = nn.Dense(self.action_dim, name="action_dense")(x) # Final Q-Values
Appendix C Experiment Details
C.1 Hyperparameters
Table 1: Atari Hyperparameters for PQN, PQN (ResNet-15) and Hadamax-PQN. These hyperparameters are equal to the original hyperparameters from the PQN baseline [17].
Parameter Value
NUM_ENVs 128
NUM_STEPS 32
EPS_START 1.0
EPS_FINISH 0.001
EPS_DECAY 0.1
NUM_EPOCHS 2
NUM_MINIBATCHES 32
NORM_INPUT False
NORM_TYPE layer_norm
LR 0.00025
MAX_GRAD_NORM 10
LR_LINEAR_DECAY False
GAMMA 0.99
LAMBDA 0.65
OPTIMIZER RAdam
C.2 Environments
We run experiments on the Atari-57 suite, where there are 57 different games in total. No per-game tuning is allowed and the same agent architecture, hyper-parameters and pre-processing needs to run on every game. The suite contains varying games that can be used to examine different properties of RL agents, e.g. long-horizon credit assignment, partial observability, hard exploration, etc.
Each observation consists of 4 grayscale images of the game state stacked together, i.e. $(4,64,64)$. The action space is discrete, and each action represents a different operation in the game. The reward function depends on the environment chosen. More details on each game can be found at https://ale.farama.org.
Atari-3 and Atari-10: We examine C51, DQN and Rainbow on Atari-3 or Atari-10 [2], which are a small but representative subset of the full Atari-57 suite. Atari-3 includes Battle Zone, Name This Game and Phoenix. Atari-10 includes Amidar, Bowling, Frostbite, Kung Fu Master, River Raid, Battle Zone, Double Dunk, Name This Game, Phoenix and Q*Bert.
C.3 Baseline Implementations
PQN: We use the official codebase  [^1] https://github.com/mttga/purejaxql of PQN and default hyper-parameter settings.
Rainbow, C51, DQN: We use implementations from cleanrl  [^2] https://github.com/vwxyzjn/cleanrl and default hyper-parameter settings.
Hadamax encoder: Since the whole PQN codebase is in Jax, we implement the Hadamax encoder for PQN in Jax as well. As Implementations of Rainbow, C51 and DQN from cleanrl are in PyTorch, we also implement the Hadamax encoder for these agents in PyTorch.
C.4 Compute Usage
We run all our experiments on a HPC cluster equipped with A100 GPUs. Each run of Hadamax-PQN needs around 45 minutes for 40 millions frames and PQN needs around 20 minutes.
Appendix D Metrics
D.1 Median Human-Normalized Score
For each game, compute the average score $x_{i}$ across multiple independent seeds. Then compute the normalized score $Z_{i}$ as:
$Z_{i}=\frac{x_{i}-r_{i}}{h_{i}-r_{i}}$
where $x_{i}$ is the raw score, and $r_{i}$ and $h_{i}$ are the random and human scores for game $i$, respectively (see Table 2 for values). After computing the normalized scores for all 57 games * seeds, they are sorted and the median value is computed.
D.2 Atari-57 Score Profile
x-axis:($\tau$ Normalized Score). Represents the threshold score (e.g., Human-Normalized Score). Higher values mean better performance.
y-axis: $\tau\%$ = fraction of games above $\tau$. Shows the fraction of games for which the agent’s normalized score is greater than $\tau$. For example, at $\tau=1$, the y-value represents what fraction of games the agent beats $\tau=1$ human performance on. In other words, it represents the percentage of games that has scores higher than $\tau$.
D.3 Atari-3 and Atari-10
The Atari-3 and Atari-10 scores approximate the median normalized score across the full 57-game Atari benchmark using subsets of 3 and 10 games, respectively [2]. The computation involves the following steps:
  1. 1.
    For each game in the subset, compute the normalized score $Z_{i}$ as:
    $Z_{i}=100\times\frac{x_{i}-r_{i}}{h_{i}-r_{i}}$
    where $x_{i}$ is the raw score, and $r_{i}$ and $h_{i}$ are the random and human scores for game $i$, respectively (see Table 2 for values).
  2. 2.
    Apply the log transform:
    $\phi(Z_{i})=\log_{10}(1+\max(0,Z_{i}))$
  3. 3.
    Compute the weighted sum $f=\sum_{i\in I}c_{i}\phi(Z_{i})$, where $I$ is the subset of games and $c_{i}$ are the subset-specific coefficients.
  4. 4.
    Obtain the predicted median score as:
    $\hat{t}=10^{f}-1$
For Atari-3, the subset comprises Battle Zone, Name This Game, and Phoenix, with coefficients $c_{i}=[0.3706,0.5133,0.1015]$.
For Atari-10, the subset includes Amidar, Bowling, Frostbite, Kung Fu Master, River Raid, Battle Zone, Double Dunk, Name This Game, Phoenix, and Q*Bert, with coefficients $c_{i}=[0.0825,0.0559,0.0691,0.0986,0.0486,0.1888,0.0852,0.1287,0.1643,0.0592]$.
Appendix E Additional Experiments
E.1 Deeper Hadamax Networks
As network scaling has become a topic of interest in the field of RL [34, 48, 41], we provide experiments using deeper versions of our encoder: 5-layer and 7-layer Hadamax-PQN. Specifically, the second and third convolutional layers in the original 3-layer encoder are duplicated, and we refrain from max-pooling the duplicates to avoid excessive compression. Similar to the ablations, the deep networks are tested on the full 57-game Atari suite for 40M environment frames. The results can be seen in Fig. 11.
Figure 11: Hadamax encoder depth Ablations.
Simply using deeper convolutional Hadamax encoders does not seem to improve performance. Although there are more promising ways to scale the Hadamax encoder both in depth and width, the computational cost was the limiting factor in pursuing this in more detail. As discussed in the main paper, we leave this as a promising research area for future work.
E.2 Hadamax with Other Agents
We modify the encoders of the widely-used cleanrl [28] implementations of C51, DQN, and Rainbow to demonstrate that the Hadamax encoder can generalize across various model-free agents. See Figure 12, on Atari-10, Hadamax improves the performance of the original C51 by $70\%$, and on Atari-3, it boosts DQN and Rainbow by $20\%$ and $30\%$, respectively. These substantial gains, achieved by simply replacing the encoder, suggest that Hadamax could serve as a new default encoder for model-free reinforcement learning methods on Atari.
Figure 12: Performance gains of DQN, C51 and Rainbow with Hadamax encoders on a subset of Atari-57.
E.3 Per-game improvement over Rainbow-DQN
Appendix F Individual Game Scores
Table 2: Final 200M frame scores.
Game Hadamax-PQN PQN (Resnet-15) PQN
Alien-v5 20045.4 16935.0 2916.3
Amidar-v5 1774.5 944.1 740.3
Assault-v5 26426.7 11160.5 15089.7
Asterix-v5 274915.0 95400.0 287697.6
Asteroids-v5 39328.0 4232.7 21047.6
Atlantis-v5 715750.0 516357.8 831884.7
BankHeist-v5 1260.2 1446.1 1336.2
BattleZone-v5 92951.2 55106.5 44130.8
BeamRider-v5 49480.9 16315.8 18131.7
Berzerk-v5 57497.4 19597.8 6061.3
Bowling-v5 29.7 30.2 42.5
Boxing-v5 99.8 96.3 98.3
Breakout-v5 607.4 470.3 489.6
Centipede-v5 17901.7 9266.5 8178.2
ChopperCommand-v5 203593.5 26974.7 11688.8
CrazyClimber-v5 202048.2 162776.7 168732.4
Defender-v5 360287.5 48761.1 66381.0
DemonAttack-v5 135450.5 125870.4 131320.0
DoubleDunk-v5 -1.8 -1.2 -1.2
Enduro-v5 2323.4 462.7 2284.6
FishingDerby-v5 46.6 31.2 45.4
Freeway-v5 33.7 21.4 33.8
Frostbite-v5 7689.5 6537.2 5623.8
Gopher-v5 67829.0 26859.5 40834.5
Gravitar-v5 1547.2 514.4 1107.3
Hero-v5 30617.4 24912.9 18099.9
IceHockey-v5 16.3 -2.4 -1.4
Jamesbond-v5 4244.8 1285.8 1942.8
Kangaroo-v5 13177.3 8728.6 13992.5
Krull-v5 10554.4 9497.9 9802.2
KungFuMaster-v5 36751.9 24102.8 38233.3
MontezumaRevenge-v5 0.0 0.0 0.0
MsPacman-v5 6968.3 4584.7 4909.7
NameThisGame-v5 21334.2 18754.3 16437.0
Phoenix-v5 267080.2 41001.4 120959.5
Pitfall-v5 -43.5 -34.4 -50.5
Pong-v5 21.0 20.8 21.0
PrivateEye-v5 3.6 3.9 7.5
Qbert-v5 25970.2 21818.4 22449.6
Riverraid-v5 29423.9 18669.8 24133.3
RoadRunner-v5 190019.6 52925.0 76600.9
Robotank-v5 71.5 66.1 68.3
Seaquest-v5 129408.8 43559.8 11554.4
Skiing-v5 -29971.3 -29479.8 -29972.3
Solaris-v5 1884.2 863.5 2189.4
SpaceInvaders-v5 22258.0 13800.3 15125.0
StarGunner-v5 549350.4 215397.7 264413.1
Surround-v5 9.4 7.5 6.3
Tennis-v5 23.8 22.9 -1.0
TimePilot-v5 17946.0 11924.0 12320.1
Tutankham-v5 258.7 216.8 248.0
UpNDown-v5 191857.2 82743.3 270833.7
Venture-v5 940.7 0.0 18.1
VideoPinball-v5 522510.3 416690.8 463022.1
WizardOfWor-v5 21526.1 13130.5 22214.2
YarsRevenge-v5 444710.8 119951.1 111611.7
Zaxxon-v5 31400.8 14229.4 17644.0
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims made in the abstract and introduction accurately reflect our paper’s contributions and scope, which is the creation of a new state-of-the-art baseline via encoder synthesis for model-free RL.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In the Conclusions and Future Work section, we dedicated a full paragraph to our view on the limitations of this work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: As this paper focuses on neural network architectures in reinforcement learning, it does not contain any theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The full, easily implementable JAX-code for the Hadamax encoder is provided in Appendix B.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: As stated in 4, we provide open access to the code of the exact convolutional architecture used for the main results in the paper.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Our experimental setup is equal to that of the baselines that are used. We specify the hyperparameters in Appendix C.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Our experiments are based on the full Atari-57 suite, which consists of 57 different pixel-based environments. Following conventional score metric visualization, we use the median-human-normalized scores over 57 environments, which are conventionally not accompanied by error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Figure 1, we show the compute times of all the algorithms used for a single 200M environment frame seed.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This paper conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This work is not tied to particular applications, and serves to optimize RL algorithms in the simulation domain. We therefore do not expect any societal impacts of this research.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our data and models used do not have a high risk for misuse. We use openly available data from the Atari-57 domain.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators of the baseline models have been properly credited in our research by way of GitHub repository links and several citations.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The code of the proposed Hadamax encoder in this paper is clearly provided in Appendix B.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Tis paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs have not been used for core components of our paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.